# Functional Protein Design with Local Domain Alignment

### v2

@@ 4.6 Ablation Study
Table 4: Ablation study of the effectiveness of each component. Performance of $\text{SR}_{1}$ for zinc-finger and immunoglobulin domains.
@@
4.7 Case study: Joint Generation with Domain and Property Annotations
Due to the flexibility of combining annotations, PAAG can jointly generate the proteins with both specific domain and property. Specifically, we generate the zinc-finger proteins in membrane-bound by the model in Section 4.5. We showcase two examples of the generated results in Figure  6. As shown in Figure  6, PAAG can successfully generate proteins under the conditions derived from both domain and property annotations, demonstrating its potential in complex protein design tasks.
Figure 6: Visualization of jointly generation with domain and property annotations. PAAG is capable of integrating domain and property annotations.
5 Related Work
@@
Moltimodal Representation Learning. By harnessing the potential of extensive image-text pair data, the Contrastive Language-Image Pretraining (CLIP) model, as proposed by Radford et al. Radford et al. (2021), employs contrastive learning to align the representations between image and text modalities.Following by CLIP, many image-text pertaining model are proposed, such as BILP Li et al. (2022), BLIP-2 Li et al. (2023) and ClipCap Mokady et al. (2021). Beyond the image-text pretraining, several studies introduce the more modalities, such as videos Xu et al. (2021), audios Tang et al. (2023) and even molecules Liu et al. (2022) into a unified representation. Specifically, for multimodal learning on protein sequences, OntoProtein Zhang et al. (2022) first learns protein representations by combining them with textual descriptions in a knowledge graph. ProtST Xu et al. (2023) constructs a large-scale dataset containing aligned pairs of protein sequences and property descriptions, and pretrain a protein-biotext model to improve performance on downstream predictive task and enables zero-shot retrieval.
@@
Protein Generation Model. With huge success of language models Devlin et al. (2019), several studies Elnaggar et al. (2021); Shin et al. (2021); Ferruz et al. (2022); Rives et al. (2019) treat protein sequences consisting chains of amino acids as a type of “languages” and pretrain models on millions of protein sequences. Upon the pretrained model, they generate the protein sequences in an autoregressive manner. In addition to the autoregressive model, Evodiff Alamdari et al. (2023b) extracts the evolutionary information from protein sequences and proposes an evolution-guided diffusion model to generate protein sequences. Given the significance of structural information for protein function, a group of methods Hsu et al. (2022); Ingraham et al. (2019); Tan et al. (2022); Jing et al. (2020), known as inverse folding, utilize the structure as a conditional input, allowing for the generation of amino acid sequences. Some studies Shi et al. (2022); Watson et al. (2023) integrate both sequential and structural information and propose a co-design model that accepts sequences and structures as conditions. Recently, ProteinDT Liu et al. (2023b) first proposes a text-sequence alignment framework, enabling its capabilities for text-guided protein generation and editing. Chroma Ingraham et al. (2023) trains a protein caption model ProCap and utilize it as a classifier guidance to generate proteins via a diffusion model.
6 Conclusion
@@
In this paper, we present PAAG, a multi-modality framework that first incorporates the rich annotation information derived from protein database, achieving the superior performance in various applications, such as representation learning and annotation-guided protein design. Crucially, we demonstrate that it is possible to use the flexible combinations of various kinds of textual annotations to guide the protein design process. We hope that PAAG will expand the possibilities of protein design and establish a robust foundation for future advancements in protein-related applications. Future research directions include functional protein editing, co-design of protein sequence-structure within alignment frameworks, and exploring the potential of larger protein dataset with lower annotation quality.
@@ Appendix A Experimental settings
A.1 General settings
Backbone Models of PAAG: Unless otherwise specified, we utilize the weights from ProtBert-BFD Elnaggar et al. (2021) as the initialization of our protein encoder. For the text encoder, we choose SciBert Beltagy et al. (2019) as the initial backbone. The scarcity of training data does not allow us to train a large decoder, we utilize a lighter decoder, initialized by DistilProtBert Geffen et al. (2022). A decoder that contains more parameters is also our future direction.
@@ A.1 General settings
A.2 ProtAnnotation and the template function $G(\cdot)$
Figure 5 depicts additional statistics of ProtAnnotation. We demonstrate several example data samples in ProtAnnotation as well as the corresponding textual description generated by the template function $G(\cdot)$.
Table 5: The statistics of ProtAnnotation
@@ A.2 ProtAnnotation and the template function $G(\cdot)$
Table 6: Examples of annotations and description of protein in ProtAnnotation
@@
Q8Y2D7 [254, 288], FPG-type Zinc finger organism_name: Pseudomonas solanacearum,
protein_name: Formamidopyrimidine-DNA glycosylase,
length: 288,
SIMILARITY: Belongs to the FPG family
This is Formamidopyrimidine-DNA glycosylase protein, from Pseudomonas solanacearum, belongs to the FPG family, contains 288 amino acids, has 1 FPG-type Zinc finger.
A.3 Biological background of zinc-finger and immunoglobulin domain
We provide some introductions of zinc-finger and immunoglobulin domain from biological perspective, elucidating their significance and functions within the context of molecular biology.
  • Zinc-finger: Zinc finger Klug and Rhodes (1987) domains are compact protein structural motif with multiple protrusions that interact with target molecules. Initially identified in the transcription factor TFIIIA from Xenopus laevis as a DNA-binding motif, these domains are now recognized for their ability to bind DNA, RNA, proteins, and lipids. Their binding characteristics depend on the amino acid sequence, the linker structure, the higher-order structures, and the number of fingers. Often found in clusters, each finger within the Zinc finger domain can have unique binding specificities. Despite being present in several unrelated protein superfamilies, these domains maintain stable structures and play a crucial role in diverse cellular processes such as gene transcription, translation, mRNA trafficking, and more. Zinc-binding motifs, part of the Zinc finger domains, rarely undergo significant conformational changes upon target binding.
  • Immunoglobulin domain: Immunoglobulin (Ig) domains Brummendorf (1995), also referred to as immunoglobulin-like domains or Ig-like domains, are protein domains that of importance in the immune system and are found in a wide range of proteins, both within and beyond the immune system. These domains are characterized by their highly conserved structure, which typically consists of 70 to 110 amino acid residues that form a compact, globular fold. The Ig domains are ubiquitous in nature and can be found in a diverse array of proteins, including antibodies, T-cell receptors, major histocompatibility complex (MHC) molecules, cell adhesion molecules, and various receptor proteins. The primary function of Ig domains is to facilitate protein-protein interactions, particularly in the context of the immune system, where they mediate the recognition, binding, and neutralization of foreign antigens.
An illustration of Zinc-finger and Immunoglobulin domains is shown in Figure 7.
Figure 7: Visualization of typical proteins containing Zinc-finger (P26634, left figure) protein and Immunoglobulin domain (A0M8Q6, right figure).
A.4 The evaluation metric for the unconditional generation
@@
A.5 More details of baselines
@@
ProGen Madani et al. (2023) first forms each property of protein as keyword tags. In the training stage, ProGen prepends the keyword tags to the corresponding amino acid sequence, and try to reconstruct the amino acid sequence in the auto-regressive manner. ProGen consequently has the ability to generate the protein sequence given the keyword tags. During controllable generation, ProGen first translates desired properties into keyword tags, then can controllably generate protein sequence.
@@
Chroma Ingraham et al. (2023) is also a protein design model that but focus more on generating the protein backbone structure. Specifically, Chroma is a diffusion model that can generate protein backbone with desired properties and functions by accepting the classifier guidance. Therefore, Chroma introduces ProCap, a protein-to-text model, as the classifier guidance to enable text-guided protein design. With an additional sequence sampling model, Chroma is capable of sampling both structure and sequence based on the conditions.
In unconditional generation, PAAG only accept the length $l$ of the protein sequence. We also define this is the $k$-th protein we generated Consequently, we set the template function as “protein number k, contains l amino acids.”
@@
While generate proteins with functional domains, we randomly sample the organism name, length and similarity from natural proteins that have corresponding functional domain. We specify the generative hyper-parameters in A.8. When given property annotations. Deeploc Almagro Armenteros et al. (2017) contains training data and test data in its original split. We utilize template function $G()$ to form the textual descriptions of test data. Then PAAG can generate protein with properties given these descriptions.
A.6 Training configurations
A.6.1 Predictive Task
In downstream prediction tasks, PAAG utilizes the embedding in the aligned space. Specifically, we not only use the protein encoder to extract the protein representations. We also keep the projector head to extract the more informative aligned features. Furthermore, we also incorporate dropout rate and weight decay to improve the generalization ability.
A.6.2 Protein Design with Domain Annotations
@@
While training PAAG with ProtAnnotation, we introduce momentum contrast He et al. (2020), and we set the queue size as 16384, and momentum as 0.995. In the contrastive learning, we initialize the learnable temperature as 0.07. Moreover, we align the our latent space in 256 dimension with one linear layer. Also, following the setting of Li et al. (2022), to enable soft labeling, we initialize the learnable alpha as 0.4. When fine-tuning on the dataset only contains zinc-finger and immunoglobulin domains, we keep these hyper-parameters but decrease queue size to 4096 since the number training data also decreases.
Besides, we also set the weight decay as 0.05 for AdamW during the learning.
A.6.3 Protein Design with Property Annotations
Due to the scarcity of training data, we conduct fine-tuning process based on the models trained in Section 4.4 for 100 epochs. We mostly keep hyper-parameters as training on ProtAnnotation but decrease the learning rate to 1e-5 without warming up.
@@ A.6.3 Protein Design with Property Annotations
For binary localization properties, soluble and membrane-bound, we set template function $G()$ as “is soluble” or “is membrane-bound”. For 10-class subcellular localization, $G()$ is “locates in {location}”.
@@
Table 7: Configurations of full-model tuning of PAAG-ProtBert on three task types. Abbr., lr.: learning rate; lrr.: learning rate ratio; dr.: dropout rate; wd.: weight decay; bs.: batch size; MSE: mean squared error; CE: cross entropy; BCE: binary cross entropy.
Task optimizer lr. lrr. dr. wd. bs. #epochs loss
@@
Bin Adam $2.0\times 10^{-5}$ 0.15 0 0 4 100 BCE
Sub Adam $2.0\times 10^{-5}$ 0.15 0.1 0 4 100 CE
@@
$\rho$-lac Adam $2.0\times 10^{-4}$ 0.02 0 $3.0\times 10^{-4}$ 36 100 MSE
AAV Adam $9.0\times 10^{-4}$ 0.02 0 0 36 100 MSE
Thermo Adam $2.0\times 10^{-4}$ 0.02 0 0 6 100 MSE
Flu Adam $2.0\times 10^{-4}$ 0.02 0.3 $4.0\times 10^{-4}$ 36 100 MSE
Sta Adam $5.0\times 10^{-4}$ 0.02 0.7 0 36 100 MSE
Table 8: Configurations of full-model tuning of PAAG-ESM-2 on three task types. Abbr., lr.: learning rate; lrr.: learning rate ratio; dr.: dropout rate; wd.: weight decay; bs.: batch size; MSE: mean squared error; CE: cross entropy; BCE: binary cross entropy.
Task optimizer lr. lrr. dr. wd. bs. #epochs loss
@@
Bin Adam $1.0\times 10^{-5}$ 0.15 0 0 4 100 BCE
Sub Adam $2.0\times 10^{-5}$ 0.15 0 0 4 100 CE
@@
$\rho$-lac Adam $1.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
AAV Adam $8.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
Thermo Adam $3.0\times 10^{-4}$ 0.02 0.6 0 2 100 MSE
Flu Adam $3.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
Sta Adam $8.0\times 10^{-5}$ 0.02 0 0 16 100 MSE
A.7 Additional experimental results
A.7.1 Results of protein generation with 10-class subcellular location annotation
We represent the confusion matrix of 10-class subcellular location annotation result in Figure 8 for clearer illustration.
Figure 8: Confusion matrix of protein design within 10 subcellular localization annotations. Numbers indicate the count of generated instances given the condition annotation in the column while predicted as row annotation.
A.8 Hyper-parameters
A.8.1 Hyper-parameters of generative task
In our generative task, we utilize nucleus sampling in our decoder. Specifically, we will sample the amino acids based on their probability instead of always choosing the amino acids with the highest probability. We set the Top-p hyperparameter as 0.9 in nucleus sampling, which means the decoder will only consider the most likely amino acids that make up 90% of the probability mass. Besides, we set the repetition penalty of decoder as 1.2.
A.8.2 Hyper-parameters of the predictive task
The hyperparameters for the predictive task concerning PAAG-ProtBert and PAAG-ESM-2 can be found in Table  7 and Table  8, respectively.

### v3

@@ 4.7. Ablation Study
Table 4. Ablation study for each component. Performance of $\text{SR}_{1}$ for zinc-finger and immunoglobulin domains.
@@
5. Related Work
@@
Moltimodal Representation Learning. By harnessing the potential of extensive image-text pair data, the Contrastive Language-Image Pretraining (CLIP) model, as proposed by Radford et al. (Radford et al., 2021), employs contrastive learning to align the representations between image and text modalities.Following by CLIP, many image-text pertaining model are proposed, such as BILP (Li et al., 2022), BLIP-2 (Li et al., 2023), InstructBLIP (Dai et al., 2023) and ClipCap (Mokady et al., 2021). Beyond the image-text pretraining, several studies introduce the more modalities, such as videos (Xu et al., 2021), audios (Tang et al., 2023) and even molecules (Liu et al., 2022; Zhang et al., 2024) into a unified representation. Specifically, for multimodal learning on protein sequences, OntoProtein (Zhang et al., 2022) first learns protein representations by combining them with textual descriptions in a knowledge graph. ProtST (Xu et al., 2023) constructs a large-scale dataset containing aligned pairs of protein sequences and property descriptions, and pretrain a protein-biotext model to improve performance on downstream predictive task and enables zero-shot retrieval.
@@
Protein Generation Model. With huge success of language models (Devlin et al., 2019), several studies (Elnaggar et al., 2021; Shin et al., 2021; Ferruz et al., 2022; Rives et al., 2019) treat protein sequences consisting chains of amino acids as a type of “languages” and pretrain models on millions of protein sequences. Upon the pretrained model, they generate the protein sequences in an autoregressive manner. In addition to the autoregressive model, Evodiff (Alamdari et al., 2023b) extracts the evolutionary information from protein sequences and proposes an evolution-guided diffusion model to generate protein sequences. Given the significance of structural information for protein function, a group of methods (Hsu et al., 2022; Ingraham et al., 2019; Tan et al., 2022; Jing et al., 2020), known as inverse folding, utilize the structure as a conditional input, allowing for the generation of amino acid sequences. Some studies (Shi et al., 2022; Watson et al., 2023) integrate both sequential and structural information and propose a co-design model that accepts sequences and structures as conditions. Recently, ProteinDT (Liu et al., 2023b) first proposes a text-sequence alignment framework, enabling its capabilities for text-guided protein generation and editing. Chroma (Ingraham et al., 2023) trains a protein caption model ProCap and utilize it as a classifier guidance to generate proteins via a diffusion model.
Although ProteinDT (Liu et al., 2023b) and ProtST (Xu et al., 2023) similarly align the protein and textual representations, the multi-level alignment framework enables PAAG to capture the both local and global properties of the protein. Furthermore, the momentum encoders and an additional matching loss assist PAAG to more effectively align the protein-level annotation with the proteins.
6. Conclusion
@@
This paper presents PAAG, a multi-modality framework that first incorporates the rich annotation information derived from protein database, achieving the superior performance in various applications, such as representation learning and annotation-guided protein design. Crucially, we demonstrate that it is possible to use the flexible combinations of various kinds of textual annotations to guide the protein design process. We hope that PAAG will expand the possibilities of protein design and establish a robust foundation for future advancements in protein-related applications. Future research directions include functional protein editing, co-design of protein sequence-structure within alignment frameworks, and exploring the potential of larger protein dataset with lower annotation quality.
Acknowledgements.
The authors would like to thank Siying Xu (China Academy of Art) for her help in visualizing the framework and results. This work was jointly supported by the following projects: the National Natural Science Foundation of China (No. 62376276); Beijing Nova Program (No. 20230484278); Beijing Outstanding Young Scientist Program (No. BJJWZYJH012019100020098); the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (23XNKJ19); Public Computing Cloud, Renmin University of China.
@@ Appendix A Experimental settings
A.1. General settings
Backbone Models of PAAG: We use ProtBert-BFD (Elnaggar et al., 2021) to initialize our protein encoder and SciBert (Beltagy et al., 2019) for the text encoder. Due to limited training data, we opt for a lighter decoder initialized by DistilProtBert (Geffen et al., 2022). Exploring a decoder with more parameters is a future direction.
@@ A.1. General settings
A.2. ProtAnnotation and template function $G(\cdot)$
Table 5 depicts additional statistics of ProtAnnotation. We demonstrate several example data samples in ProtAnnotation as well as the corresponding textual description generated by the template function $G(\cdot)$.
Table 5. The statistics of ProtAnnotation
@@ A.2. ProtAnnotation and template function $G(\cdot)$
Table 6. Examples of annotations and description of protein
in ProtAnnotation
@@
Table 7. The success rate on the prompt with single domain.
Method SR1(Zinc finger) SR1(Immunoglobulin)
Chroma 0.33% 0%
ProteinDT 1.67% 22%
PAAG 22.67% 51%
PAAGsingle ${}_{\text{single domain}}$ 18.33% 46%
Table 8. Results of the proteins unconditionally generated by fine-tuned PAAG.
Model Distinc-2 Diversity Novelty
Natural 0.4309(0) 0.829(0) -
ProGen 0.3003(30.31%) 0.845(1.93%) 0.374
Chroma 0.3211(25.48%) 0.855(3.14%) 0.638
PAAGbefore ${}_{\text{before ft}}$ 0.4314(0.12%) 0.815(1.69%) 0.766
PAAGafter ${}_{\text{after ft}}$ 0.4524(4.99%) 0.822(0.84%) 0.674
A.3. Biological background of zinc-finger and immunoglobulin domain
We here introduce zinc-finger and immunoglobulin domains, highlighting their biological significance and functions in biology.
  • Zinc-finger: Zinc finger (Klug and Rhodes, 1987) domains are compact protein motifs known for binding DNA, RNA, proteins, and lipids, with binding characteristics influenced by amino acid sequence, linker structure, and number of fingers. Despite appearing in various protein families, they maintain stable structures and are crucial in processes like gene transcription and mRNA trafficking.
  • Immunoglobulin domain: Immunoglobulin (Ig) domains play a key role in protein-protein interactions, especially within the immune system, where they help recognize, bind, and neutralize antigens. These domains have a conserved structure found in a variety of proteins, including antibodies and cell adhesion molecules.
Examples of Zinc-finger and Ig domains are shown in Figure 7.
Figure 7. Visualization of typical proteins containing Zinc-finger (P26634) and Immunoglobulin (A0M8Q6).
A.4. The evaluation metric for the unconditional generation
@@
A.5. More details of baselines
@@
ProGen (Madani et al., 2023) represents each protein property as keyword tags, which are prepended to amino acid sequences during training to enable auto-regressive reconstruction. This allows ProGen to generate protein sequences from keyword tags. For controllable generation, ProGen translates desired properties into keyword tags to guide sequence generation.
@@
Chroma (Ingraham et al., 2023) is a protein design model focusing on generating protein backbones. It uses classifier guidance and ProCap, a protein-to-text model, for text-guided designs. With a sequence sampling model, Chroma can generate both structure and sequence based on specified conditions.
ProteinDT (Liu et al., 2023b) is a text-guided protein design model that aligns protein and text representations, then uses aligned text representations to design protein sequences via autoregressive and diffusion models. For fair comparison, we use autoregressive ProteinDT as our baseline. In unconditional generation, due to the lack of a specific template in ProteinDT, we randomly sample prompts from its dataset to evaluate the quality of generated proteins. In conditional experiments, ProteinDT receives the same prompts as PAAG.
@@
In unconditional generation, PAAG accepts only the protein sequence length $l$ and the order $k$ of protein generated, using the template: ”protein number k, contains l amino acids.” For generating proteins with functional domains, we randomly sample organism names, lengths, and similarities from natural proteins with the corresponding domains. Generative hyper-parameters are specified in A.6. Using property annotations, we apply the template function $G()$ to describe test data from the Deeploc (Almagro Armenteros et al., 2017) dataset, enabling PAAG to generate proteins with specified properties.
A.6. Training configurations
A.6.1. Predictive Task
In downstream prediction tasks, PAAG uses embeddings in the aligned space by employing both the protein encoder and the projector head for more informative aligned features. Hyperparameters for predictive tasks using PAAG-ProtBert and PAAG-ESM-2 are detailed in Table  9 and Table  10.
A.6.2. Protein Design with Domain Annotations
@@
While training PAAG with ProtAnnotation, we incorporate momentum contrast (He et al., 2020) with a queue size of 16,384 and momentum of 0.995. The learnable temperature in contrastive learning is set to 0.07, and our latent space is aligned to 256 dimensions with a linear layer. Following (Li et al., 2022), we initialize learnable alpha at 0.4 for soft labeling. When fine-tuning on datasets with only zinc-finger and immunoglobulin domains, we maintain these hyperparameters but reduce the queue size to 4,096 due to less training data. Additionally, weight decay for AdamW is set at 0.05.
In our generative task, we use nucleus sampling in the decoder, sampling amino acids based on probability rather than always selecting the highest probability ones. The Top-p hyperparameter is set to 0.9, meaning the decoder considers the top 90% probability mass. We also set the decoder’s repetition penalty to 1.2.
A.6.3. Protein Design with Property Annotations
Due to limited training data, we fine-tune the models from Section 4.4 for 100 epochs. We retain the ProtAnnotation training hyperparameters but lower the learning rate to 1e-5 without warm-up.
@@ A.6.3. Protein Design with Property Annotations
For binary localization properties, soluble and membrane-bound, we set template function $G()$ as “is soluble” or “is membrane-bound”.
@@
Table 9. Configurations of full-model tuning of PAAG-ProtBert. Abbr., lr.: learning rate; lrr.: learning rate ratio; dr.: dropout rate; wd.: weight decay; bs.: batch size; MSE: mean squared error; CE: cross entropy; BCE: binary cross entropy.
Task lr. lrr. dr. wd. bs. #epochs loss
@@
Bin $2.0\times 10^{-5}$ 0.15 0 0 4 100 BCE
Sub $2.0\times 10^{-5}$ 0.15 0.1 0 4 100 CE
@@
$\rho$-lac $2.0\times 10^{-4}$ 0.02 0 $3.0\times 10^{-4}$ 36 100 MSE
AAV $9.0\times 10^{-4}$ 0.02 0 0 36 100 MSE
Thermo $2.0\times 10^{-4}$ 0.02 0 0 6 100 MSE
Flu $2.0\times 10^{-4}$ 0.02 0.3 $4.0\times 10^{-4}$ 36 100 MSE
Sta $5.0\times 10^{-4}$ 0.02 0.7 0 36 100 MSE
Table 10. Hyperparameters for PAAG-ESM-2 in downstream tasks.
Task lr. lrr. dr. wd. bs. #epochs loss
@@
Bin $1.0\times 10^{-5}$ 0.15 0 0 4 100 BCE
Sub $2.0\times 10^{-5}$ 0.15 0 0 4 100 CE
@@
$\rho$-lac $1.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
AAV $8.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
Thermo $3.0\times 10^{-4}$ 0.02 0.6 0 2 100 MSE
Flu $3.0\times 10^{-4}$ 0.02 0 0 16 100 MSE
Sta $8.0\times 10^{-5}$ 0.02 0 0 16 100 MSE
A.7. Additional experimental results
A.7.1. Results of generating proteins with EGF-like domain
We here generates $300$ functional proteins with EGF-like domain for ProGen, ProteinDT and PAAG. The results are in Table 11.
Table 11. Success Rate of generating proteins with EGF-like domain.
EFG-like domain SR100 SR10 SR1 SR0.1 SR0.01 SR0.001
ProGen 1% 1% 1% 1% 1% 1%
ProteinDT 0% 0% 0% 0% 0% 0%
PAAG 28.67% 28.33% 26.00% 22.33% 17.33% 11.67%
A.7.2. More evaluation on the proteins generated by fine-tuned PAAG
To ensure fine-tuned PAAG isn’t just memorizing the training set, we evaluate distinct-2, diversity, and novelty for unconditionally generated sequences. Table 8 shows that, even after fine-tuning, PAAG produces proteins with higher novelty than Chroma and ProGen, indicating it doesn’t overfit, thanks to label smoothing. Additionally, fine-tuned PAAG’s diversity is closest to natural proteins, suggesting it captures similar amino acid distributions.